### v2

@@ 4 Mode evolution
$\Omega^{2}(t)=e^{-\alpha\phi}m_{\chi}^{2}-a^{-2}\nabla^{2}+\xi_{\chi}R\\[5.69046pt] +3\alpha\left(\dfrac{1}{6}-\xi_{\chi}\right)(\ddot{\phi}+3H\dot{\phi})-\dfrac{3\alpha^{2}}{2}\left(\dfrac{1}{6}-\xi_{\chi}\right)\dot{\phi}^{2}\,.$ (34)
@@
have positive frequency $\omega_{k}^{2}(t)>0$ even if $m_{\mathrm{eff}}^{2}<0$, which corresponds to sub-horizon regime $k\gg aH$. This occurs when the modes are well inside the horizon at the beginning of inflation. In the figure 3 we show the evolution of $\omega_{k}^{2}(t)$ for a non-minimal coupling $\xi=0$. Before $t_{\mathrm{cross}}$, the frequency is dominated by the decay of the physical mode $k/a\sim e^{-Ht}$. Once the mode crosses the horizon $(k\ll aH)$, $\omega_{k}^{2}(t)$ remains almost constant, and its value is negative since the coefficient of the dominant term $H^{2}$ is negative. For $t\gtrsim t_{\mathrm{end}}$, the frequency increases and oscillates after inflation as the background terms oscillate around the minimum of the potential (preheating).
@@ 5 Description of particle production
This behavior usually happens at early times $t_{i}\to 0$ and at sufficiently late times $t_{f}\to+\infty$, but for intermediate times $t_{i}<t_{*}<t_{f}$ during the early universe, the evolution of the modes is no longer adiabatic $|A_{k}|\gg 1$. During inflation, the positive-frequency solution corresponds to the Bunch-Davies vacuum (43), as described in the previous section. This evolution remains valid as long as the parameter $|A_{k}|\ll 1$ at all times. While this condition is met, no particle production occurs, as adiabatic evolution is ensured, and there is no mixing of negative- and positive-frequency modes. From figure 3, we can see that the evolution of $\omega_{k}(t)$ is smooth until the end of inflation. At that point, the condition $|A_{k}|\ll 1$ is strongly violated, leading to an enhancement in particle production, and gravitational particle production is enhanced at the time $t_{*}$, corresponding to the end of inflation. We show the non-adiabaticity evolution in the figure 4 parametrized by the adiabatic parameter $|A_{k}(t)|$. The mixture of positive and negative frequency modes at time $t_{\mathrm{end}}$ can be represented as the sum of the solutions $\exp(\pm i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime})$ using the WKB approximation [14, 5]
$X_{k}(t)=\dfrac{\alpha(t)}{\sqrt{2\omega_{k}(t)}}\,e^{-i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}+\dfrac{\beta(t)}{\sqrt{2\omega_{k}(t)}}\,e^{i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}\,,$
where $\alpha_{k}(t)$ and $\beta_{k}(t)$ are the time-dependent Bogolyubov coefficients. Comparing with the Bunch-Davies initial condition (45), one see thats $\alpha_{k}(t_{0})=1$ and $\beta_{k}(t_{0})=0$.
In other words, gravitational particle production occurs at very early times $t_{\mathrm{in}}\to 0$ and at late times $t_{f}\to\infty$ if the initial and final vacuum states are not the same. We define the initial vacuum state with the corresponding creation-annihilation operators $\hat{a}^{\dagger}_{\mathbf{k}}$ and $\hat{a}_{\mathbf{k}}$ by
@@
The physical number density $n_{\chi}(t)$ in the late-time limit $t\to\infty$ is given by
$a^{3}(t)\,n_{\chi}(t)=\int_{0}^{\infty}\mathrm{d}\log k\,n_{k}(t)\,,\quad n_{k}(t)=\dfrac{k^{3}|\beta_{k}(t)|^{2}}{2\pi^{2}}\,,$
where $n_{k}(t)$ is the comoving number density of particles with comoving wavenumber $k$. Notice that the expression (49) is only valid when $\omega_{k}^{2}(t)>0$, so the comoving number density of particles created is valid at times when the frequency (39) is positive.
@@ 5.1 Numerical results
What follows is to find the spectrum of $\chi$ particles and their evolution in the cosmic time $t$. Hence we numerically solve (38), constructed the Bogolyubov coefficient $|\beta_{k}(t)|$, and compute the spectrum $n_{k}$ using (50). In the previous section, we described the procedure to numerically evaluate the set of equations to ensure that each mode $k$ is within the horizon at the beginning of the evolution, which allows us to use the Bunch-Davies vacuum (43) as an initial condition of the mode equation for $X_{K}(t)$. The initial conditions for the background are set at time $t_{\mathrm{ini}}$, corresponding to the moment when the modes are within the horizon. For example, for $N_{*}=60$ at the time of crossing, the initial time corresponds to $t_{\mathrm{ini}}\approx 28.712$ in units of $[10^{-5}M_{p}^{-1}]$, and the corresponding mode $k=aH$ will be the value of $aH$ at time $t_{\mathrm{ini}}$. Then, the Bogolyubov coefficient $|\beta_{k}(t)|^{2}$ is calculated using (49), and the comoving number density of gravitational particle production is determined by $n_{k}(t)$.
@@ 5.1.1 Conformal coupling $\xi_{\chi}=1/6$
We first present the results for conformally coupled particles ($\xi_{\chi}=1/6$). As an example, we present the time evolution of the comoving number of particles created $n_{k}(t)$ as a function of time for two different cases: light particles $m_{\chi}<H_{\mathrm{end}}$ and heavy particles $m_{\chi}\gtrsim H_{\mathrm{end}}$. We should note that the expression for $\beta_{k}(t)$ is only valid when $\omega_{k}^{2}(t)>0$ , we point out that $n_{k}(t)$ only makes physical sense at late times $t\to\infty$ [5]. The frequency (39) for scalar conformally coupled turns out to be
@@
which can be negative for $k\ll aH$. For this reason, we take into account the sub-horizon evolution of the modes $X_{k}(t)$ and we take $k$ deep inside the horizon at the beginning of the simulation. In the figure, we present two cases: $m_{\chi}=10^{-5}H_{\mathrm{end}}$ and $m_{\chi}=10^{-1}H_{\mathrm{end}}$. In both panels of figure 5, we observe that the number of particles created saturates at the end of inflation, stabilizing and becoming constant at late times $t\to\infty$. This growth in $n_{k}$ is due to the non-adiabaticity of $\omega_{k}$ at the end of inflation, as can be seen in figure 4.
Figure 5: Comoving number density of particles, plotted on a logarithmic scale, as a function of cosmic time $t$ for two different values of the spectator mass $m_{\chi}$, with the same mode $k$, corresponding to $N_{k}=60$. The graph depicts the evolution of the occupation number $n_{k}/k_{\mathrm{end}}^{3}$, where $k_{\mathrm{end}}=a_{\mathrm{end}}H_{\mathrm{end}}$ denotes the mode at the end of inflation
The oscillations of $n_{k}(t)$ after inflation arise from resonances occurring at specific moments between the oscillations of the modes $X_{k}$ and the frequency itself. During this epoch, particle production is driven primarily by the oscillatory behavior of the scalaron and the Higgs field, as illustrated in figure 2, through the term $\dot{H}$ and the Ricci scalar $R$ via the non-minimal coupling $\xi_{\chi}=1/6$. Additionally, we observe that particle production near the end of inflation is most efficient for light particles. In both cases, the particle number is well-defined, grows from the end of inflation, and stabilizes at late times, as expected.
@@
We have numerically computed the spectra of the comoving number density $n_{k}(t)$ as a function of $k/k_{\mathrm{end}}$ for a range of light masses varying from $10^{-5}H_{\mathrm{end}}$ to $10^{-1}H_{\mathrm{end}}$. Our findings are summarized in figure 6. From the figure, one can observe several remarkable features. Conformally coupled scalars show a spectra that are strongly peaked at some value $k_{\mathrm{peak}}$. For IR modes $(k\ll k_{\mathrm{end}})$, we see that the spectrum is blue-tilted, reaching a peak at some intermediate value $k_{\mathrm{peak}}>k_{\mathrm{end}}$, and then is suppressed for UV modes $(k\gg k_{\mathrm{end}})$ due to inflation, that is, those that remain within the horizon during inflation. IR modes leave the horizon during inflation and re-enter later during the reheating epoch after inflation. Therefore, the dynamics of the $X_{k}(t)$ modes are dictated by the effective mass term (41) with a quasi-De Sitter spacetime as background with $H\simeq\mathrm{constant}$. It is important to note that the UV modes do not cross the horizon during inflation. In place, these modes are mainly produced during the transition to the matter-dominated era and in general during the reheating era. The suppression of UV modes occurs because the $\chi$ scalars become supermassive at the end of inflation due to the exponentially scalaron-dependence mass, causing the effective mass to become substantial. As a result, gravitational particle production becomes less efficient, leading to a significant suppression in the particle density number $n_{k}(t)$. The mass of the field plays a key role in determining the amplitude of $n_{k}$ at the peak momentum $k_{\mathrm{peak}}$, which systematically decreases as $m_{\chi}$ increases [43].
For large masses $m_{\chi}\gtrsim H_{\mathrm{end}}$, the situation is quite similar. As the mass increases, the particle production tends to decrease due to the increase of the effective mass. This can be seen in the left panel of figure 7 for a mass of $m_{\chi}=H_{\mathrm{end}}$ and for a mode crossing the horizon at $N_{k}=60$. In the rigt panel, we present the comoving spectra of particle production for $m_{\chi}\gtrsim H_{\mathrm{end}}$. We find that the behaviour for UV modes is different than in the case of light masses. The location of the peak $k_{\mathrm{peak}}$ is different as a function of the mass value, once $k_{\mathrm{end}}$. While $k_{\mathrm{peak}}>k_{\mathrm{end}}$ as in the case of light particles, the entry to the horizon, and hence the tail suppression of the UV modes is different. This is largely due to the possible resonances that exist for large masses, as the mass term $e^{-\alpha\phi}$ dominates after inflation when $\phi<M_{p}$, which can excite different modes after inflation.

### v3

@@ 4 Mode evolution
$\Omega^{2}(t)=e^{-\alpha\phi}m_{\chi}^{2}+\dfrac{k^{2}}{a^{2}}+\xi_{\chi}R+3\alpha\left(\dfrac{1}{6}-\xi_{\chi}\right)(\ddot{\phi}+3H\dot{\phi})-\dfrac{3\alpha^{2}}{2}\left(\dfrac{1}{6}-\xi_{\chi}\right)\dot{\phi}^{2}\,.$ (29)
@@
have positive frequency $\omega_{k}^{2}(t)>0$ even if $m_{\mathrm{eff}}^{2}<0$, which corresponds to sub-horizon regime $k\gg aH$. This occurs when the modes are well inside the horizon at the beginning of inflation. In the Fig. 3 we show the evolution of $\omega_{k}^{2}(t)$ for a non-minimal coupling $\xi=0$. Before $t_{\mathrm{cross}}$, the frequency is dominated by the decay of the physical mode $k/a\sim e^{-Ht}$. Once the mode crosses the horizon $(k\ll aH)$, $\omega_{k}^{2}(t)$ remains almost constant, and its value is negative since the coefficient of the dominant term $H^{2}$ is negative. For $t\gtrsim t_{\mathrm{end}}$, the frequency increases and oscillates after inflation as the background terms oscillate around the minimum of the potential (preheating).
@@ 5 Description of particle production
This behavior usually happens at early times $t_{i}\to 0$ and at sufficiently late times $t_{f}\to+\infty$, but for intermediate times $t_{i}<t_{*}<t_{f}$ during the early universe, the evolution of the modes is no longer adiabatic $|A_{k}|\gg 1$. During inflation, the positive-frequency solution corresponds to the Bunch-Davies vacuum (38), as described in the previous section. This evolution remains valid as long as the parameter $|A_{k}|\ll 1$ at all times. While this condition is met, no particle production occurs, as adiabatic evolution is ensured, and there is no mixing of negative- and positive-frequency modes. From Fig. 3, we can see that the evolution of $\omega_{k}(t)$ is smooth until the end of inflation. At that point, the condition $|A_{k}|\ll 1$ is strongly violated, leading to an enhancement in particle production, and gravitational particle production is enhanced at the time $t_{*}$, corresponding to the end of inflation $t_{\mathrm{end}}$. We show the non-adiabaticity evolution in Fig. 4 parametrized by the adiabatic parameter $|A_{k}(t)|$. The mixture of positive and negative frequency modes at time $t_{\mathrm{end}}$ can be represented as the sum of the solutions $\exp(\pm i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime})$ using the WKB approximation [14, 5]
$X_{k}(t)=\dfrac{\alpha_{k}(t)}{\sqrt{2\omega_{k}(t)}}\,e^{-i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}+\dfrac{\beta_{k}(t)}{\sqrt{2\omega_{k}(t)}}\,e^{i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}\,,$ (41)
where $\alpha_{k}(t)$ and $\beta_{k}(t)$ are the time-dependent Bogolyubov coefficients. Comparing with the Bunch-Davies initial condition (40), one see thats $\alpha_{k}(t_{0})=1$ and $\beta_{k}(t_{0})=0$. Gravitational particle production is a quantum process that occurs in the early universe, primarily near the end of inflation and continues for a few e-folds afterwards until the $\beta$-Bogolyubov coefficient is stabilized, allowing for a well-defined computation of the energy density of the produced particles. However, if the initial and final vacuum states are not the same, particle production can also be interpreted in terms of asymptotic in and out states, highlighting its connection to both early and late-time behavior. We define the initial vacuum state with the corresponding creation-annihilation operators $\hat{a}^{\dagger}_{\mathbf{k}}$ and $\hat{a}_{\mathbf{k}}$ by
@@
The total number comoving density $n_{\chi}(t)$ of created particles in the late-time limit $t\to\infty$ is given by
$a^{3}(t)\,n_{\chi}(t)=\int_{0}^{\infty}\dfrac{\mathrm{d}^{3}k}{(2\pi)^{3}}|\beta_{k}(t)|^{2}\,,$ (45)
which can be rewritten as
$a^{3}(t)n_{\chi}(t)=\int\dfrac{\mathrm{d}k}{k}n_{k}(t)\,,\quad n_{k}(t)=\dfrac{k^{3}}{2\pi^{2}}|\beta_{k}(t)|^{2}$ (46)
where $n_{k}(t)$ is the comoving spectral number density of particles. Notice that the expression (44) is only valid when $\omega_{k}^{2}(t)>0$, so the comoving number density of particles created is valid at times when the frequency (34) is positive. In some cases, it is possible to analytically evaluate the $\beta$-Bogolyubov coefficient. The expression (44) is typically used in numerical approaches, as, in most cases, evaluating $\beta_{k}$ numerically requires knowledge of the mode dynamics $X_{k}(t)$ and the frequency $\omega_{k}(t)$. That is, we need to numerically solve the equation (33) with Bunch-Davies initial conditions. However, there is an alternative method to compute the Bogolyubov coefficients $\alpha_{k}$ and $\beta_{k}$. The WKB approximation (41) provides a solution to (33) if the time-dependent Bogolyubov coefficients $\alpha_{k}(t)$ and $\beta_{k}(t)$ satisfy the system of equations given in [14].
$\dot{\alpha}_{k}=\dfrac{\dot{\omega}_{k}}{2\omega_{k}}e^{2i\psi_{k}(t)}\,\beta_{k}\,,\quad\dot{\beta}_{k}=\dfrac{\dot{\omega}_{k}}{2\omega_{k}}e^{-2i\psi_{k}(t)}\,\alpha_{k}\,,$ (47)
where $\psi_{k}(t)=\int_{t_{i}}^{t}\mathrm{d}t^{\prime}\,\omega_{k}(t^{\prime})$ is a time-dependent oscillatory phase. Gravitational particle production occurs in the Bunch-Davies regime, where $|\beta_{k}|\ll 1$ and $|\alpha_{k}|\simeq 1$. In this case, the Bogolyubov coefficient $\beta_{k}$ is given by
$|\beta_{k}(t)|\simeq\dfrac{1}{2}\int_{t_{i}}^{t}\mathrm{d}t^{\prime}\,\dfrac{\dot{\omega}_{k}}{\omega_{k}}\exp\left(-2i\int_{t_{i}}^{t^{\prime}}\mathrm{d}t^{\prime\prime}\omega_{k}(t^{\prime\prime})\right)\,.$ (48)
It is possible to numerically solve this integral without the need to numerically solve the equation (33) for an initial time $t_{i}\ll t$; however, the numerical calculation of the integral becomes complex in most cases when the oscillatory phase $\psi_{k}(t)$ oscillates rapidly. This approach is more useful when one seeks to evaluate $\beta_{k}$ analytically or approximately. To do so, the steepest descent method is commonly employed in the literature [10, 12, 64, 68]. We have chosen to adopt a numerical approach to compute (45), due to the nature of the model, which involves a non-minimal coupling and a nontrivial potential term. These features significantly complicate the evaluation of the integral expression (48). As this analysis lies beyond the main scope of the present work, we leave it for future investigations.
@@ 5.1 Numerical results
What follows is to find the spectrum of $\chi$ particles and their evolution in the cosmic time $t$. Hence we numerically solve (33), constructed the Bogolyubov coefficient $|\beta_{k}(t)|$, and compute the spectrum $n_{k}$ using (45). In the previous section, we described the procedure to numerically evaluate the set of equations to ensure that each mode $k$ is within the horizon at the beginning of the evolution, which allows us to use the Bunch-Davies vacuum (38) as an initial condition of the mode equation for $X_{k}(t)$. The initial conditions for the background are set at time $t_{\mathrm{ini}}$, corresponding to the moment when the modes are within the horizon. For example, for $N_{*}=60$ at the time of crossing, the initial time corresponds to $t_{\mathrm{ini}}\approx 28.712$ in units of $[10^{-5}M_{p}^{-1}]$, and the corresponding mode $k=aH$ will be the value of $aH$ at time $t_{\mathrm{ini}}$. In this way, we can associate a value of $k$ to each value of $N_{e}$, which determines the scale $k$ that crosses the horizon $N_{e}$ e-folds before the end of inflation.
Figure 5: The Bogolyubov coefficient $\beta_{k}$ is shown as a function of cosmic time $t$ for a spectator field with masses $m_{\chi}=10^{-1}H_{\mathrm{end}}$ and $m_{\chi}=3H_{\mathrm{end}}$, corresponding to a mode that exits the horizon $N_{k}=60$ e-folds before the end of inflation. We observe that $\beta_{k}$ increases near the end of inflation and eventually stabilizes at late times, for $t\gg t_{\mathrm{end}}$.
An alternative approach to selecting the $k$-modes is to work directly with the dimensionless ratio $k/k_{\mathrm{end}}$ in the equation for $X_{k}(t)$. In this framework, modes with $k<k_{\mathrm{end}}$ will have crossed the horizon at some point before the end of inflation. However, for our purposes, it is more convenient to generate a list of modes that satisfy $k=aH$ by solving the background equations and associating each mode with the corresponding number of e-folds $N_{e}$ before the end of inflation. The resulting list of physical modes $k$ is then normalized by $k_{\mathrm{end}}$ to construct the spectrum $n_{k}(t)$. The Bogolyubov coefficient $|\beta_{k}(t)|^{2}$ is computed using Eq. (44), and the comoving number density of gravitationally produced particles is determined by $n_{k}(t)$. It is important to note that the particle number is evaluated only once it has stabilized; that is, we extract the value of $n_{k}$ at the end of the numerical simulation to ensure that it has reached a steady state.
@@ 5.1.1 Conformal coupling $\xi_{\chi}=1/6$
We first present the results for conformally coupled particles ($\xi_{\chi}=1/6$). As an example, we present the time evolution of the $\beta$-Bogolyubov coefficient as a function of time for two different cases: light particles $m_{\chi}<H_{\mathrm{end}}$ and heavy particles $m_{\chi}\gtrsim H_{\mathrm{end}}$. We should note that the expression for $\beta_{k}(t)$ is only valid when $\omega_{k}^{2}(t)>0$ , we point out that $n_{k}(t)$ only makes physical sense at late times $t\to\infty$ [5] when its value stabilizes.
The frequency (34) for scalar conformally coupled turns out to be
@@
which can be negative for $k\ll aH$. For this reason, we take into account the sub-horizon evolution of the modes $X_{k}(t)$ and we take $k$ deep inside the horizon at the beginning of the simulation. We present the two different cases in the Fig. 5: $m_{\chi}=10^{-1}H_{\mathrm{end}}$ and $m_{\chi}=3H_{\mathrm{end}}$. One can observe that the number of particles created saturates at the end of inflation, stabilizing and becoming constant at late times $t\to\infty$. This growth in $\beta_{k}(t)$ is due to the non-adiabaticity of $\omega_{k}$ at the end of inflation, as can be seen in Fig. 4. The oscillations of $\beta_{k}(t)$ after inflation arise from resonances occurring at specific moments between the oscillations of the modes $X_{k}$ and the frequency itself. During this epoch, particle production is driven primarily by the oscillatory behavior of the scalaron and the Higgs field, as illustrated in Fig. 2, through the term $\dot{H}$ and the Ricci scalar $R$ via the non-minimal coupling $\xi_{\chi}=1/6$. For both cases, the particle number is well-defined, grows from the end of inflation, and stabilizes at late times, as expected. We can also observe that particle production is more efficient for lighter species, as indicated by the value of $|\beta_{k}(t)|^{2}$ once particle production has ceased. In particular, we find $|\beta_{k}(t_{\mathrm{final}})|^{2}\approx 0.0071$ for the case $m_{\chi}=10^{-1}H_{\mathrm{end}}$, while for $m_{\chi}=3H_{\mathrm{end}}$ we obtain $|\beta_{k}(t_{\mathrm{final}})|^{2}\approx 1.216\times 10^{-6}$. This behavior is precisely what one expects: heavier particles are produced less efficiently via gravitational mechanisms, as more energy is required to excite them from the vacuum.
@@
We have also numerically computed the spectra of the comoving number density $n_{k}(t)$ as a function of $k/k_{\mathrm{end}}$ for a range of light masses varying from $10^{-5}H_{\mathrm{end}}$ to $10^{-1}H_{\mathrm{end}}$. Our findings are summarized in Fig. 6. From the figure, one can observe several remarkable features. Conformally coupled scalars show a spectra that are strongly peaked at some value $k_{\mathrm{peak}}$. For IR modes $(k\ll k_{\mathrm{end}})$, we see that the spectrum is blue-tilted, reaching a peak at some intermediate value $k_{\mathrm{peak}}>k_{\mathrm{end}}$, and then is suppressed for UV modes $(k\gg k_{\mathrm{end}})$ due to inflation, that is, those that remain within the horizon during inflation. IR modes leave the horizon during inflation and re-enter later during the reheating epoch after inflation. Therefore, the dynamics of the $X_{k}(t)$ modes are dictated by the effective mass term (36) with a quasi-De Sitter spacetime as background with $H\simeq\mathrm{constant}$. This results in a comoving number density with a scaling $n_{k}\sim k^{3}$ for these modes. That is, in the IR region, the particle occupation number vanishes $n_{k}\to 0$ as $k\to 0$, consistent with what is observed in the figure. On the other hand, UV modes correspond to those that never crossed the horizon during inflation, and therefore their production is strongly suppressed. In place, these modes are mainly produced during the transition to the matter-dominated era and in general during the reheating era. This behavior is characteristic of gravitational particle production and has been extensively reported in the literature [69, 70, 49, 48, 50].
The field’s mass $m_{\chi}$ plays a key role in determining the amplitude of $n_{k}$ at the peak momentum $k_{\mathrm{peak}}$, which systematically decreases as $m_{\chi}$ increases [48, 50, 51]. For large masses $m_{\chi}\gtrsim H_{\mathrm{end}}$, the situation is quite similar. As the mass increases, the particle production tends to decrease due to the increase of the effective mass. This behavior can be clearly seen in the particle spectrum $n_{k}$ shown in Fig. 7.
@@
The case with conformal coupling $\xi_{\chi}=1/6$ exhibits a similar behavior to the light mass case $m_{\chi}\ll H_{\mathrm{end}}$: in the infrared (IR) region, the spectrum scales as $n_{k}\sim k^{3}$, while in the ultraviolet (UV) region, particle production is exponentially suppressed.